# RIS-Empowered OTFS Modulation With Faster-than-Nyquist Signaling in High-Mobility Wireless Communications

### v1

@@ Abstract
High-mobility wireless communication systems suffer from severe Doppler spread and multi-path delay, which degrade the reliability and spectral efficiency of conventional modulation schemes. Orthogonal time frequency space (OTFS) modulation offers strong robustness in such environments by representing symbols in the delay-Doppler (DD) domain, while faster-than-Nyquist (FTN) signaling can further enhance spectral efficiency through intentional symbol packing. Meanwhile, reconfigurable intelligent surfaces (RIS) provide a promising means to improve link quality via passive beamforming. Motivated by these advantages, we propose a novel RIS-empowered OTFS modulation with FTN signaling (RIS-OTFS-FTN) scheme. First, we establish a unified DD-domain input–output relationship that jointly accounts for RIS passive beamforming, FTN-induced inter-symbol interference, and DD-domain channel characteristics. Based on this model, we provide comprehensive analytical performance for the frame error rate, spectral efficiency, and peak-to-average power ratio (PAPR), etc. Furthermore, a practical RIS phase adjustment strategy with quantized phase selection is designed to maximize the effective channel gain. Extensive Monte Carlo simulations under a standardized extended vehicular A (EVA) channel model validate the theoretical results and provide key insights into the trade-offs among spectral efficiency, PAPR, input back-off (IBO), and bit/frame error performance, with some interesting insights. The proposed RIS-OTFS-FTN scheme demonstrates notable performance gains in both reliability and spectral efficiency, offering a viable solution for future high-mobility and spectrum-constrained wireless systems.
@@ I-C Reconfigurable Intelligent Surface
Thus, a novel and popular wireless technique called the Reconfigurable intelligent surfaces (RIS) is introduced in our works. RIS has recently emerged as a promising enabling technology for beyond fifth-generation (B5G) and sixth-generation (6G) wireless communication systems. An RIS typically consists of a two-dimensional (2D) planar array of reconfigurable reflection elements (REs), each fabricated from engineered materials capable of manipulating incident electromagnetic waves in a programmable manner [15]. Each RE can be individually controlled by an intelligent controller or chip, which receives configuration instructions from the transmitter and adjusts the phase shifts of the reflected signals accordingly [16]. This unique capability enables RIS to reshape the wireless propagation environment by steering the reflected beams toward desired directions, thereby extending coverage, enhancing received signal power, and improving link reliability. Unlike conventional active relays such as unmanned aerial vehicles or vehicular relays, RIS operates without radio-frequency (RF) chains, power amplifiers, or active signal regeneration [17]. As a result, it can achieve high beamforming gains with minimal power consumption, a feature commonly referred to as passive beamforming [18]-[20]. Through intelligent phase shift optimization, RIS can enhance the signal-to-noise ratio (SNR) at the receiver while incurring negligible additional energy cost, making it highly attractive for energy-efficient wireless communications. Since FTN causes degradation in bit-error rate (BER) performance and OTFS still fails to achieve sufficiently satisfactory performance in high-mobility scenarios, RIS can be considered to address these issues. In our work, RIS provides benefits beyond simple SNR enhancement. When FTN-aided interference dominates, merely increasing transmit power is inefficient and limited by power amplifiers (PA) back-off (IBO) and peak-to-average power ratio (PAPR)-induced nonlinearity. By configuring its phases, RIS can reshape the cascaded DD-domain channel to improve post-equalization reliability under FTN. This reduces the required transmit power and expands the feasible IBO region for a target AFER and SE, demonstrating a non-trivial synergy between RIS and OTFS-FTN beyond a mere combination of two.
@@ I-D Related Works
Recently, several studies investigate the RIS-assisted OTFS (RIS-OTFS) and FTN-assisted OTFS (FTN-OTFS) schemes. On the one hand, the pure RIS-OTFS schemes are first proposed and analyzed in [21]-[25], where RIS significantly improves the BER performance of OTFS systems. In addition, fundamental aspects such as the input–output relationship, RIS phase adjustment design, multiple-input multiple-output (MIMO) system modeling, and channel estimation are thoroughly discussed in these works. However, further theoretical investigation of RIS-OTFS is still required, particularly in terms of error performance, spectral efficiency, and the power fluctuation characteristics of modulated signals, in order to provide deeper insights for practical applications. Later, [26] proposes a novel index modulation-aided RIS-OTFS scheme to enhance spectral efficiency, but it further increases the detection complexity that is already high in OTFS system. As another high-efficiency technology, FTN serves as a better alternative to index modulation, offering lower detection complexity while supporting higher data rates. On the other hand, various FTN-OTFS schemes are proposed and examined in [27]-[30]. While these schemes improve the spectral efficiency of OTFS systems, they often suffer from degraded BER performance. In particular, when frequency compression is applied, as in [29]-[30], the BER performance further deteriorates due to the increased ISI. Overall, both of these fields lack sufficient literature to provide a stronger theoretical foundation and more insights, leaving a research gap.
@@ I-E Contributions
OTFS mitigates Doppler spread in high-mobility scenarios but is limited to Nyquist-rate SE. FTN surpasses this limit via compressed symbol spacing, yet introduces severe ISI that degrades reliability. To address this, we employ RIS. Its passive beamforming gain effectively offsets the SNR penalty incurred by FTN. This combined approach supports aggressive FTN compression, unlocking high SE without compromising the reliability of OTFS in challenging channels, thereby overcoming the traditional mobility-SE trade-off. In light of the above works and motivated aforementioned challenges in reliability and spectral efficiency, a novel wireless scheme called the RIS-empowered OTFS modulation with FTN signaling (RIS-OTFS-FTN) scheme is proposed and investigated in this paper. Main contributions of our works are briefly summarized as
@@
  • A tailored RIS phase adjustment strategy that optimizes the discrete phase shifts of individual reflection elements based on the DD channel state information is designed. This design aligns the reflected signal components constructively at the receiver, thereby maximizing the effective channel gain. The proposed method accounts for practical hardware constraints by employing quantized phase control, while still achieving near-optimal performance.
@@ I-F Organization and Notation
Organization: The rest of the paper is organized as follows: Section II presents the system model of the proposed RIS-OTFS-FTN scheme, including the signal model in the DD domain and the corresponding input–output relationship. Section III provides the theoretical performance analysis, covering the AFER, spectral efficiency, and PAPR, etc. An optimal strategy of RIS phase adjustment is designed in Section IV. Section V presents extensive simulation results under the standardized EVA channel model to validate the analytical findings and to provide additional performance insights. Finally, Section VI concludes the paper.
@@ II-C Received Signals
where $\beta_{q}e^{j\theta_{q}}$ is the reflection coefficience of $q$-th reflection element with amplitude attenuation $\beta_{q}\in\left[0,1\right]$ and adjusted phase $\theta_{q}\in\left[0,2\pi\right]$. Next, we can simply obtain the received signals from $q$-th to receiver, which is given as
@@ II-D Input-Output Relation in DD Domain
where $\tilde{\mathbf{s}}\in\mathbb{C}^{NM\times 1}$ and $\tilde{\mathbf{y}}\in\mathbb{C}^{NM\times 1}$ respectively represents the transmitted and the received data vector in time domain with $\tilde{\mathbf{y}}=\left\{y\left[a\right]\right\}_{a=0}^{MN-1}$ and $\tilde{\mathbf{s}}=\left\{s\left[a\right]\right\}_{a=0}^{MN-1}$, and $\tilde{\mathbf{H}}_{q}\in\mathbb{C}^{NM\times 1}$ is the cascaded channel matrix with cascaded coefficience $h_{q}$.
@@
Next, applying SFFT as Eq. (15), the matrix-form system output in the DD domain becomes [27]:
$\mathbf{y}=\mathbf{H}^{\mathrm{eff}}\mathbf{X}^{\mathrm{DD}}+\mathbf{z}_{\mathrm{DD}},$ (21)
where $\mathbf{H}^{\text{eff}}=[\mathbf{H}_{1},\mathbf{H}_{2},\dots,\mathbf{H}_{Q}]\,\boldsymbol{\Phi}$ with $\mathbf{\Phi}=\mathrm{diag}\left(\beta_{1}e^{j\theta_{1}},\dots,\beta_{Q}e^{j\theta_{Q}}\right)\otimes\mathbb{I}_{NM}$ and $\mathbf{H}_{q}\in\mathbb{C}^{NM\times NM}$ is the channel coefficient matrices of the cascaded channel, and $\mathbf{z}_{\mathrm{DD}}=\left(\mathbf{F}_{N}\otimes\mathbf{F}_{M}^{H}\right)\mathbf{z}_{\mathrm{TF}}$, $\mathbf{z}_{\mathrm{TF}}\in\mathbb{C}^{MN\times 1}$ is the vector of time domain additive complex Gaussian noise with covariance $\mathbb{E}\left[\mathbf{z}_{\mathrm{TF}}{\mathbf{z}_{\mathrm{TF}}}^{H}\right]=\mathbf{\Sigma}_{\mathbf{z}}=\sigma^{2}\mathbf{G}$, $\mathbf{G}\in\mathbb{R}^{MN\times MN}$ is a Toeplitz matrix and $\left\{\mathbf{G}\right\}_{\boldsymbol{n},\boldsymbol{m}}=g[n-m]\triangleq g((n-m)\alpha T)$.
@@ II-E Linear Minimum Mean Square Error Detector
In this work, taking into account the trade-off between computational complexity and error performance, the Linear Minimum Mean Square Error (LMMSE) detector is employed for the proposed scheme, instead of the maximum likelihood (ML) or message-passing (MP) detectors. Although the vehicle-RIS channel matrix $\mathbf{H}_{1}$ and RIS-user channel matrix $\mathbf{H}_{2}$ are individually sparse, the cascaded effective channel matrix $\mathbf{H}^{\mathrm{eff}}=\mathbf{H}_{2}\mathbf{\Phi}\mathbf{H}_{1}$ loses its sparsity after matrix multiplication [23]. This leads to increased complexity and degraded error performance for the MP detector, as the density of the factor graph increases significantly. Second, the optimal ML detector requires an exhaustive search over all potential symbol combinations, calculating Euclidean distances for the entire $NM$-dimensional vector space. This results in prohibitively high computational complexity that scales exponentially with the grid size, rendering it impractical for high-mobility OTFS systems. Consequently, we adopt the LMMSE detector [31], which minimizes the mean square error under the constraint of a linear filter. The LMMSE detection is given by
$\hat{\mathbf{x}}^{\mathrm{DD}}=\left(\mathbf{H}_{\mathrm{eff}}^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}+\mathbf{I}_{NM}\right)^{-1}\mathbf{H}_{\mathrm{eff}}^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{y},$ (22)
where $\hat{\mathbf{x}}^{\mathrm{DD}}\in\mathbb{C}^{NM\times 1}$ denotes the estimated vectorized signal in the DD domain, and $\mathbf{H}_{\mathrm{eff}}^{\dagger}$ is the conjugate transpose of $\mathbf{H}^{\mathrm{eff}}$. Here, $\mathbf{\Sigma}_{z}=\sigma^{2}\mathbf{G}$ represents the covariance matrix of the colored noise, where $\mathbf{G}$ captures the correlation introduced by the non-orthogonal FTN pulse shaping. It is worth noting that the term $\mathbf{\Sigma}_{z}^{-1}$ implicitly performs noise whitening, ensuring optimality within the class of linear estimators.
@@ III-A AFER
While the ML detector is challenging to implement in both practical systems and simulations, its theoretical performance can still be characterized by deriving the upper bound of the AFER, which allows us to gain valuable insights into its fundamental behavior. We define $E_{f}$ as the energy per frame with $E_{f}=MNE_{x}$. Let $\mathbf{x}^{\mathrm{DD}}=\mathrm{vec}(\mathbf{X}^{\mathrm{DD}})\in\mathbb{C}^{NM\times 1}$ denote the vectorized transmitted symbol block. According to the system model, $\mathbf{x}^{\mathrm{DD}}$ is drawn from a finite codebook $\mathcal{S}$ with size $|\mathcal{S}|$, where $\mathcal{S}=\{\mathbf{x}_{1}^{\mathrm{DD}},\mathbf{x}_{2}^{\mathrm{DD}},...,\mathbf{x}_{|\mathcal{S}|}^{\mathrm{DD}}\}$. Given that the FTN-induced noise $\mathbf{z}$ follows a colored Gaussian distribution with covariance $\mathbf{\Sigma}_{z}$, the optimal ML detector minimizes the Mahalanobis distance rather than the Euclidean distance. Thus, the detection rule is expressed as
$\hat{\mathbf{x}}^{\mathrm{DD}}=\mathrm{arg}\min_{\mathbf{x}^{\mathrm{DD}}\in\mathcal{S}}\left\|\mathbf{y}-\mathbf{H}^{\mathrm{eff}}\mathbf{x}^{\mathrm{DD}}\right\|_{\mathbf{\Sigma}_{z}^{-1}}^{2},$ (23)
@@
where $P(\mathbf{x}_{i}^{\mathrm{DD}}\rightarrow\mathbf{x}_{j}^{\mathrm{DD}})$ is the pairwise error probability (PEP). Substituting $\mathbf{y}=\mathbf{H}^{\mathrm{eff}}\mathbf{x}^{\mathrm{DD}}_{i}+\mathbf{z}$ into (23), the PEP event corresponds to the inequality:
@@
$2\Re\left\{\mathbf{z}^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}\Delta_{ij}\right\}\leq-\Delta_{ij}^{\dagger}\mathbf{H}_{\mathrm{eff}}^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}\Delta_{ij},$ (26)
where $\Delta_{ij}=\mathbf{x}^{\mathrm{DD}}_{i}-\mathbf{x}^{\mathrm{DD}}_{j}$. The term on the left-hand side, denoted as $w=\mathbf{z}^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}\Delta_{ij}$, is a scalar Gaussian random variable with zero mean. Its variance is derived as
@@
$=\Delta_{ij}^{\dagger}\mathbf{H}_{\mathrm{eff}}^{\dagger}\mathbf{\Sigma}_{z}^{-1}\underbrace{\mathbb{E}[\mathbf{z}\mathbf{z}^{\dagger}]}_{\mathbf{\Sigma}_{z}}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}\Delta_{ij}$
$=\Delta_{ij}^{\dagger}\mathbf{H}_{\mathrm{eff}}^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}\Delta_{ij}.$ (27)
@@
$P(\mathbf{x}_{i}^{\mathrm{DD}}\rightarrow\mathbf{x}_{j}^{\mathrm{DD}})$ $=Q\left(\sqrt{\frac{\Delta_{ij}^{\dagger}\mathbf{H}_{\mathrm{eff}}^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}\Delta_{ij}}{2}}\right)$  (28)
$\leq\exp\left(-\frac{\Delta_{ij}^{\dagger}\mathbf{H}_{\mathrm{eff}}^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}\Delta_{ij}}{4}\right).$

### v2

@@ Abstract
Faster-than-Nyquist (FTN) signaling offers a promising path to enhance the spectral efficiency (SE) of orthogonal time frequency space (OTFS) modulation in high-mobility scenarios. However, the inherent non-orthogonal pulse shaping induces severe structured interference and colored noise, leading to reliability degradation and error floors. This paper proposes a reconfigurable intelligent surface (RIS)-empowered OTFS-FTN framework, demonstrating that the RIS acts as a critical enabler rather than a generic SNR booster. By strategically configuring RIS phases, we reshape the cascaded delay-Doppler (DD) channel to mitigate FTN-induced impairments, thereby enhancing reliability and relaxing power amplifier (PA) back-off constraints. We establish a unified DD-domain model explicitly accounting for FTN-induced colored noise and derive analytical expressions for the average frame error rate (AFER) upper bound, SE, peak-to-average power ratio (PAPR), and input back-off (IBO). Moreover, a low-complexity RIS phase design is proposed, with quantified computational and control overheads. Extensive simulations against comprehensive baselines, including Nyquist OTFS and power-compensated benchmarks, validate that the proposed framework achieves a superior SE-reliability-IBO trade-off. Results show up to around 10 dB reduction in required SNR for equivalent SE and AFER, confirming gains that cannot be replicated by simply optimizing simpler systems.
@@ I-C Reconfigurable Intelligent Surface
Thus, a novel and popular wireless technique called the Reconfigurable intelligent surfaces (RIS) is introduced in our work. RIS has recently emerged as a promising enabling technology for beyond fifth-generation (B5G) and sixth-generation (6G) wireless communication systems. An RIS typically consists of a two-dimensional (2D) planar array of reconfigurable reflection elements each fabricated from engineered materials capable of manipulating incident electromagnetic waves in a programmable manner [16]. Each reflection element can be individually controlled by an intelligent controller or chip, which receives configuration instructions from the transmitter and adjusts the phase shifts of the reflected signals accordingly [17]. This unique capability enables RIS to reshape the wireless propagation environment by steering the reflected beams toward desired directions, thereby extending coverage, enhancing received signal power, and improving link reliability. Unlike conventional active relays such as unmanned aerial vehicles or vehicular relays, RIS operates without radio-frequency (RF) chains, power amplifiers, or active signal regeneration [18]. As a result, it can achieve high beamforming gains with minimal power consumption, a feature commonly referred to as passive beamforming [19]-[20]. Through intelligent phase shift optimization, RIS can enhance the signal-to-noise ratio (SNR) at the receiver while incurring negligible additional energy cost, making it highly attractive for energy-efficient wireless communications. Since FTN causes degradation in bit-error rate (BER) performance and OTFS still fails to achieve sufficiently satisfactory performance in high-mobility scenarios, RIS can be considered to address these issues. Despite its robustness, RIS-OTFS faces a fundamental SE ceiling imposed by the Nyquist criterion. Although high-order modulations can improve SE, they degrade severely in harsh fading channels. Consequently, existing systems suffer from a rigid robustness-efficiency trade-off, preventing the simultaneous achievement of high data rates and reliable transmission. In our work, RIS provides benefits beyond simple SNR enhancement. When FTN-aided interference dominates, only increasing transmit power is inefficient and limited by power amplifiers (PA) back-off (IBO) and peak-to-average power ratio (PAPR)-induced nonlinearity. By configuring phases, RIS can reshape the cascaded channel to improve post-equalization reliability under FTN. This reduces the required transmit power and expands the feasible IBO region for a target AFER and SE, demonstrating a non-trivial synergy between RIS and OTFS-FTN beyond a mere combination of two.
@@ I-D Related Works
Recently, several studies investigate the RIS-assisted OTFS (RIS-OTFS) and FTN-assisted OTFS (FTN-OTFS) schemes. On the one hand, the pure RIS-OTFS schemes are first proposed and analyzed in [21]-[25], where RIS significantly improves the BER performance of OTFS systems. In addition, fundamental aspects such as the input–output relationship, RIS phase adjustment design, multiple-input multiple-output (MIMO) system modeling, and channel estimation are thoroughly discussed in these works. However, further theoretical investigation of RIS-OTFS is still required, particularly in terms of error performance, spectral efficiency, and the power fluctuation characteristics of modulated signals, in order to provide deeper insights for practical applications. Later, [26] proposes a novel index modulation-aided RIS-OTFS scheme to enhance spectral efficiency, but it further increases the detection complexity that is already high in OTFS system. As another high-efficiency technology, FTN serves as a better alternative to index modulation, offering lower detection complexity while supporting higher data rates. On the other hand, various FTN-OTFS schemes are proposed and examined in [27]-[30]. While these schemes improve the spectral efficiency of OTFS systems, they often suffer from degraded BER performance. In particular, when frequency compression is applied, as in [29]-[30], the BER performance further deteriorates due to the increased ISI. Plus, [31] demonstrates a low-complexity transceiver for OTFS-FTN that applies Reverse Cuthill-McKee reordering and Cholesky factorization to the sparse channel matrix. This reduces complexity from cubic to near-linear in frame size, providing a feasible path to deploy the proposed system at scale. Overall, both of these fields lack sufficient literature to provide a stronger theoretical foundation and more insights, leaving a research gap and potential improvment among the RIS, FTN, and OTFS.
@@ I-E Contributions
OTFS mitigates Doppler spread in high-mobility scenarios but is limited to Nyquist-rate SE. FTN surpasses this limit via compressed symbol spacing, yet introduces severe ISI that degrades reliability. To address this, we employ RIS. Its passive beamforming gain effectively offsets the SNR penalty incurred by FTN. This synergy supports aggressive FTN compression, unlocking high SE without compromising the reliability of OTFS in challenging channels, thereby overcoming the traditional mobility-SE trade-off. In light of the above works and motivated aforementioned challenges in reliability and spectral efficiency, a novel wireless scheme called the RIS-empowered OTFS modulation with FTN signaling (RIS-OTFS-FTN) scheme is proposed and investigated in this paper. Main contributions of our work are briefly summarized as
@@
  • A tailored RIS phase adjustment strategy that optimizes the discrete phase shifts of individual reflection elements based on the DD channel state information is designed. This design aligns reflected signal components constructively at receiver, thus maximizing the effective channel gain. The proposed method accounts for practical hardware constraints by employing quantized phase control, while still achieving near-optimal performance.
@@ I-F Organization and Notation
Organization: Section II presents the system model of the proposed RIS-OTFS-FTN scheme, including the signal model in the DD domain and the corresponding input–output relationship. Section III provides the theoretical performance analysis, covering the AFER, spectral efficiency, and PAPR, etc. An optimal strategy of RIS phase adjustment is designed in Section IV. Section V presents extensive simulation results under the standardized EVA channel model to validate the analytical findings and to provide additional performance insights. Finally, Section VI concludes the paper.
@@ II-C Received Signals
where $\beta_{q}e^{j\theta_{q}}$ is the reflection coefficient of $q$-th reflection element with amplitude attenuation $\beta_{q}\in\left[0,1\right]$ and adjusted phase $\theta_{q}\in\left[0,2\pi\right]$. Next, we can simply obtain the received signals from $q$-th to receiver, which is given as
@@ II-D Input-Output Relation in DD Domain
where $\tilde{\mathbf{s}}\in\mathbb{C}^{NM\times 1}$ and $\tilde{\mathbf{y}}\in\mathbb{C}^{NM\times 1}$ respectively represents the transmitted and the received data vector in time domain with $\tilde{\mathbf{y}}=\left\{y\left[a\right]\right\}_{a=0}^{MN-1}$ and $\tilde{\mathbf{s}}=\left\{s\left[a\right]\right\}_{a=0}^{MN-1}$, and $\tilde{\mathbf{H}}_{q}\in\mathbb{C}^{NM\times NM}$ is the cascaded channel matrix with cascaded coefficient $h_{q}$.
@@
Let $\mathbf{X}^{\mathrm{DD}}\in\mathbb{C}^{M\times N}$ denote the transmitted symbol matrix in the DD domain, whose $(m,n)$-th entry carries the data symbol mapped onto the DD grid. For notational convenience, we define the vectorized DD-domain symbol block as $\mathbf{x}^{\mathrm{DD}}\triangleq\mathrm{vec}\!\left(\mathbf{X}^{\mathrm{DD}}\right)\in\mathbb{C}^{MN\times 1}$. Next, applying SFFT as Eq. (15), the matrix-form system output in the DD domain becomes [27]:
$\mathbf{y}=\mathbf{H}^{\mathrm{eff}}\,\mathbf{x}^{\mathrm{DD}}+\mathbf{z}^{\mathrm{DD}},$ (21)
where $\mathbf{H}^{\text{eff}}=[\mathbf{H}_{1},\mathbf{H}_{2},\dots,\mathbf{H}_{Q}]\,\boldsymbol{\Phi}$ with $\mathbf{\Phi}=\mathrm{diag}\left(\beta_{1}e^{j\theta_{1}},\dots,\beta_{Q}e^{j\theta_{Q}}\right)\otimes\mathbb{I}_{NM}$ and $\mathbf{H}_{q}\in\mathbb{C}^{NM\times NM}$ is the channel coefficient matrices of the cascaded channel  [^2] Specifically, each element of the effective channel matrix $\mathbf{H}_{q}\in\mathbb{C}^{NM\times NM}$ at row $u$ and column $v$ (where $0\leq u,v\leq NM-1$) is explicitly given by the discretized ambiguity function as $\mathbf{H}_{q}]_{u,v}=\sum_{p=1}^{L}h_{p}e^{j\vartheta_{p}}\mathcal{A}_{g}\left(\frac{u-v}{NM}T_{f}-\tau_{p},\frac{u-v}{NM}\Delta f-\nu_{p}\right)$, where $L$ is the number of propagation paths, $h_{p}$ is the path gain, $\vartheta_{p}$ includes the RIS phase shift, $(\tau_{p},\nu_{p})$ are the delay and Doppler shifts of the $p$-th path and $\mathcal{A}_{g}(\tau,\nu)=\int g_{\text{tx}}(t)g_{\text{rx}}^{*}(t-\tau)e^{-j2\pi\nu t}dt$ represents the cross-ambiguity function of the transmit and receive pulse shaping filters, which inherently accounts for the inter-symbol interference induced by FTN signaling., and $\mathbf{z}_{\mathrm{DD}}=\left(\mathbf{F}_{N}\otimes\mathbf{F}_{M}^{H}\right)\mathbf{z}_{\mathrm{TF}}$, $\mathbf{z}_{\mathrm{TF}}\in\mathbb{C}^{MN\times 1}$ is the vector of time domain additive colored complex Gaussian noise with covariance $\mathbb{E}[\mathbf{z}_{\mathrm{TF}}\mathbf{z}_{\mathrm{TF}}^{H}]=\sigma^{2}\mathbf{G}$, where $\mathbf{G}$ is the Toeplitz correlation matrix determined by the FTN pulse shaping. Moreover, $\mathbf{F}_{N}\in\mathbb{C}^{N\times N}$ and $\mathbf{F}_{M}\in\mathbb{C}^{M\times M}$ denote the normalized Discrete Fourier Transform (DFT) matrices. The $(k,l)$-th entry of $\mathbf{F}_{N}$ is given by $\mathbf{F}_{N}]_{k,l}=\frac{1}{\sqrt{N}}e^{-j\frac{2\pi}{N}(k-1)(l-1)},\quad 1\leq k,l\leq N$, and $\mathbf{F}_{M}$ is defined similarly with dimension $M$.
We note that FTN signaling disrupts the standard OTFS block-circulant property, resulting in a block-Toeplitz effective channel $\mathbf{H}^{\mathrm{eff}}$. Although this precludes circular convolution-based equalization, the matrix remains sparse and banded due to localized pulse shaping, which is efficiently handled by the adopted LMMSE detector.
@@ II-E Linear Minimum Mean Square Error Detector
In this work, taking into account the trade-off between complexity and error performance, the Linear Minimum Mean Square Error (LMMSE) detector is employed for the proposed scheme, instead of the maximum likelihood (ML) or message-passing (MP) detectors. Although the vehicle-RIS channel matrix $\mathbf{H}_{1}$ and RIS-user channel matrix $\mathbf{H}_{2}$ are individually sparse, the cascaded effective channel matrix $\mathbf{H}^{\mathrm{eff}}=\mathbf{H}_{2}\mathbf{\Phi}\mathbf{H}_{1}$ loses its sparsity after matrix multiplication [23]. This leads to increased complexity and degraded error performance for the MP detector, as the density of the factor graph increases significantly. Second, the optimal ML detector requires an exhaustive search over all potential symbol combinations, calculating Euclidean distances for the entire $NM$-dimensional vector space. This results in prohibitively high computational complexity that scales exponentially with the grid size, rendering it impractical for high-mobility OTFS systems. Consequently, we adopt the LMMSE detector [32], which minimizes the mean square error under the constraint of a linear filter. The LMMSE detection based on the structured colored noise is given by:
$\hat{\mathbf{x}}^{\mathrm{DD}}=\left(\mathbf{H}_{\mathrm{eff}}^{\dagger}\mathbf{H}_{\mathrm{eff}}+\sigma^{2}\mathbf{G}\right)^{-1}\mathbf{H}_{\mathrm{eff}}^{\dagger}\mathbf{y},$ (22)
where $\hat{\mathbf{x}}^{\mathrm{DD}}\in\mathbb{C}^{NM\times 1}$ denotes the estimated signal vector. Here, $\mathbf{H}_{\mathrm{eff}}^{\dagger}$ is the conjugate transpose of the effective channel matrix $\mathbf{H}_{\mathrm{eff}}$.
@@ III-A AFER
While the ML detector is challenging to implement in both practical systems and simulations, its theoretical performance can still be characterized by deriving the upper bound of the AFER, which allows us to gain valuable insights into its fundamental behavior.  [^3] Our analysis uses the optimal ML detector as a benchmark. The derived ML upper bound rigorously captures the system’s inherent diversity order and coding gain, defining its performance potential. The LMMSE simulations, meanwhile, demonstrate the practical performance achievable under specific receiver constraints. We define $E_{f}$ as the energy per frame with $E_{f}=MNE_{x}$. Let $\mathbf{x}^{\mathrm{DD}}=\mathrm{vec}(\mathbf{X}^{\mathrm{DD}})\in\mathbb{C}^{NM\times 1}$ denote the vectorized transmitted symbol block. According to the system model, $\mathbf{x}^{\mathrm{DD}}$ is drawn from a finite codebook $\mathcal{S}$ with size $|\mathcal{S}|$, where $\mathcal{S}=\{\mathbf{x}_{1}^{\mathrm{DD}},\mathbf{x}_{2}^{\mathrm{DD}},...,\mathbf{x}_{|\mathcal{S}|}^{\mathrm{DD}}\}$.
Given that the FTN-induced noise $\mathbf{z}$ follows a colored Gaussian distribution with covariance $\mathbf{\Sigma}_{z}$, the optimal ML detector minimizes the Mahalanobis distance rather than the Euclidean distance. Thus, the detection rule is expressed as
$\hat{\mathbf{x}}^{\mathrm{DD}}$ $=\mathrm{arg}\min_{\mathbf{x}^{\mathrm{DD}}\in\mathcal{S}}(\mathbf{y}-\mathbf{H}^{\mathrm{eff}}\mathbf{x})^{H}\mathbf{\Sigma}_{z}^{-1}(\mathbf{y}-\mathbf{H}^{\mathrm{eff}}\mathbf{x})$
$=\mathrm{arg}\min_{\mathbf{x}^{\mathrm{DD}}\in\mathcal{S}}\left\|\mathbf{y}-\mathbf{H}^{\mathrm{eff}}\mathbf{x}^{\mathrm{DD}}\right\|_{\mathbf{\Sigma}_{z}^{-1}}^{2},$ (23)
@@
where $P(\mathbf{x}_{i}^{\mathrm{DD}}\rightarrow\mathbf{x}_{j}^{\mathrm{DD}})$ is the pairwise error probability (PEP). Substituting $\mathbf{y}=\mathbf{H}^{\mathrm{eff}}\mathbf{x}^{\mathrm{DD}}_{i}+\mathbf{z}$ into (III-A), the PEP event corresponds to the inequality:
@@
$2\Re\left\{\mathbf{z}^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}\Delta_{ij}\right\}\leq-\Delta_{ij}^{\dagger}(\mathbf{H}^{\mathrm{eff}})^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}\Delta_{ij},$ (26)
where $\Delta_{ij}=\mathbf{x}^{\mathrm{DD}}_{i}-\mathbf{x}^{\mathrm{DD}}_{j}$ is the error difference vector. The term on the left-hand side, denoted as $w=\mathbf{z}^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}\Delta_{ij}$, is a scalar Gaussian random variable with zero mean. Its variance is explicitly derived as
@@
$=\Delta_{ij}^{\dagger}(\mathbf{H}^{\mathrm{eff}})^{\dagger}\mathbf{\Sigma}_{z}^{-1}\underbrace{\mathbb{E}[\mathbf{z}\mathbf{z}^{\dagger}]}_{\mathbf{\Sigma}_{z}}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}\Delta_{ij}$
$=\Delta_{ij}^{\dagger}(\mathbf{H}^{\mathrm{eff}})^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}\Delta_{ij}.$ (27)
@@
$P(\mathbf{x}_{i}^{\mathrm{DD}}\rightarrow\mathbf{x}_{j}^{\mathrm{DD}})$ $=Q\left(\sqrt{\frac{\Delta_{ij}^{\dagger}(\mathbf{H}^{\mathrm{eff}})^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}\Delta_{ij}}{2}}\right)$  (28)
$\leq\exp\left(-\frac{\Delta_{ij}^{\dagger}(\mathbf{H}^{\mathrm{eff}})^{\dagger}\mathbf{\Sigma}_{z}^{-1}\mathbf{H}^{\mathrm{eff}}\Delta_{ij}}{4}\right).$